Jan 18,2000 day shift


Subject: Jan 18,2000 day shift
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:30:15 CST


              Shift Summary for day shift, Jan 18,2000

On Shift:Mike Arenton, Ricado Barbosa

Hours of beam delivered at all intensities: 7.75
Hours of beam delivered at nominal intensity: 7.75
Hours of KAMI test data: 7

Accounting of lost time: run changes, MUBEND fridge problem.

========================================================================
Runs ended this shift:
Run 15585 KAMI_LAMPPI
    15587 "
    15588 "
    15589 "
    15590 "
    15591 "
    15592 KAMI_LAMPPI including accidentals

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
  Took data using KAMI_LAMPPI to get protons from lambda's
to calibrate BAH energy scale., Got about 14K in several runs.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15585 07:00 first try at protons from lambda's
15587 07:15 change NM4AN to positive polarity, rate of
                       protons in BAH much better with this polarity.
15588 07:55 good KAMI_LAMPPI run, with lots of KFREND's
                       to fix stream 5 problems.
15589 11:30 try to add accidentals to readout,
             12:30 Abort, abort.. tape drive on ktev3 has
                       "use cleaning tape" light on.
             12:30 beam off, MUBEND.s fridge problems.
15590 13:00 new run with new tapes., no tapes ready to go,
                       had to label some.
             13:15 beam returns
15591 14:15 still another new run trying to add accidentals
                       but RK forgot to tell us one more thing that
                       needed doing.
15592 14:50 finally a run with accidentals!
             15:30 end run, ask for access to remove regenerator
                       in preparation for KE3 run.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift
  We may need more tapes to be labled.



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:30:17 CST